A runner jogs 12 km north then turns and runs 16 km east in three hours. What is his displacement? Calculate his average speed. Calculate his average velocity (including the direction).


1
Expert's answer
2021-11-14T17:13:39-0500



From the right triangle above the displacement is:


"x = \sqrt{12^2 + 16^2} = 20km"

The average speed is total distance divided by total time:


"speed = \dfrac{12km + 16km}{3h} \approx 9.3km/h"

The average velocity is the displacement divided by total time:


"velocity = \dfrac{20km}{3h} \approx 6.7km/h"


The direction of the velocity is the same as the direction of the displacement "x" (see figure).


Answer. Displacement: 20km, speed: 9.3km/h, velocity: 6.7km/h, in the direction of the displacement.